Rob is 21 years old, angie is twice as old as Rob was when Angie was old as Rob is now, how old is angie?
Found 3 solutions by greenestamps, jorel1380, stanbon:
Answer by greenestamps(13200)



Let x be the number of years ago when Angie was as old as Rob is now. Then...

Rob's age now is 21; his age x years ago was 21-x.
Angie's age now is 2(21-x) = 42-2x. (she is twice as old as Rob was x years ago)
So Angie's age x years ago was 42-3x.

But Angie's age then is equal to Rob's age now:
42-3x+=+21
21+=+3x
x+=+7

Angie's current age is 42-2x = 42-14 = 28.

Answer by jorel1380(3719)

Let n be the number of years that have passed since Angie was Rob's age, or 21. Then:
21+n=2(21-n)
21+n=42-2n
3n=21
n=7
Angie is 28, Rob is 21. Seven years ago, Rob was 14---Angie is now twice that, or 28
☺☺☺☺

Answer by stanbon(75887)

Rob is 21 years old, angie is twice as old as Rob was when Angie was old as Rob is now, how old is angie?
-------------------
Assume the constant difference in their ages is "x".
A = 2(21-x) where A-R = x
-----
So, A = 42-2x
A = 42 - 2(A-R)
A = 42 - 2A + 2R
3A = 42+2(21) = 84
A = 28 yrs (Angie's age now)
-------------------
Cheers,
Stan H.
--------